What is the probability of a face card given that the card is black?
It is 3/13. The fact that the card is black makes no difference since the probability is the same for both colours.
Pr(Red or Blue) = 0.5833...Pr(R on first pull and Blue on second pull) = 0.0533...
4
What is he probability that the sum of a pair of dice when rolled is six?
If the dice are fair then it is 5/36.
What do we call the table used to calculate the probability of an offsprings genotype?
It is called a punnet square.
The answer is 1/169.
What is a example of cluster sampling?
Cluster sampling is a scheme which is used when sampling from the whole population would be too expensive - in terms of time or money. The sample space is divided into clusters, a random sample of clusters is selected and then every member of the population within the selected clusters is studied.
For example, suppose you wanted to collect information from schools across a country and had calculated that a 5% sample was required. Rather than criss-crossing the country, you could divide the country into units: for example counties. You then select counties (your clusters) so that they cover 5% of the nation's schools. Visit each chosen county and sample all schools in it. The selection of counties would probably also need to be controlled so that urban and rural areas are properly represented.
Which of the following is the best definition of the sample space Space of a probability event?
Since there is only one definition given, it has to be better than the ones which are not even featured!
What is the probability of getting a least one tail when tossing eight fair coins?
The probability of getting at least 1 tails is (1 - probability of getting all heads)
The probability of getting all heads (no tails) is ½ x ½ x ½ x ½ x ½ x ½ x ½ x ½ = 1/256 = 0.00390625 so the probability of getting at least ONE tails is 1-0.30390625 = 0.99609375 = 255/256
There is no information on the population from which the person is selected. But if, in a total population of N people, T drive trucks, E exercise four or more times per week, and TE are truck drivers who exercise four or more times per week, then the probability is
T/N + E/N - TE/N.
What is the relationship between risk and probability?
Risk is defined as "A possibility of incurring loss or misfortune". So that means it doesn't always happen, which implies that it will happen a certain percentage of the time. This percentage is also known as the probability of it occurring, which means if you know the probability of the risk happening you can do a lot of things with it, such as find the expected value of the risk :)
If you were just told that it was possible that your house would burn down tomorrow, you wouldn't know what to expect, because you don't know the probability of this happening, but if you're told that something will happen with a 50% or even an 80% probability, you'll take the event more seriously.
What is the probability of rolling a pair of dice and obtaining a total score of 9 or greater?
It is 5/18.
If you roll a fair six sided number cube. What is the probability of rolling and eight?
The probability is 0.
The probability will depend on how much you know and the extent of guessing.
You randomly select one card from a 52-card deck. Find the probability of selecting the king of diamonds or the jack of
The theoretical probability of a given event is 9 13 explain what each number represents?
The naive answer is that the random experiment (or trial) can be modelled theoretically. This model has 13 equally likely outcomes and that 9 of them are favourable to the event. That is, 9 out of 13 equally likely outcomes involve the event occurring.
A slightly more sophisticated explanation is that there are a multiple of 13 equally likely outcomes and the same multiple of 9 are favourable to the event under study. For example, the probability of randomly drawing a non-honour from a regular deck of cards. The total number of possible outcomes is 52 (= 13*4) and the number of favourable outcomes is 36 (= 9*4). The probability, 36/52 has then been simplified to 9/13.
This answer still assumes that each individual outcome is equally likely but that is not necessary. The experiment could involve trying to hitting one of two targets: a 2*2 and a 3*3 (in the same square units of length) targets. If anywhere in the two targets has the same probability, then the theoretical probability of hitting the bigger target is 9/(9+4) = 9/13.
How many ways can you rearrange average?
This is how you do it, there are 7 letters in average so it would be, 7x6x5x4x3x2x1.